OpenStudy (liizzyliizz):

given the equation y= 3sin ^2 (x/2) what is an equation of the tangent line to the graph x=pi . ( I've done these in the past but I can't remember where to start, can someone just help me out with the steps)

OpenStudy (mertsj):

Find the y coordinate of the point.

OpenStudy (mertsj):

Then find the slope of the tangent which is the value of the first derivative when x is replaced with pi

OpenStudy (mertsj):

After you know the slope and the point, you can write the equation of the tangent line.

OpenStudy (mertsj):

First derivative is:\[y'=\frac{3\sin (x)}{2}\]

OpenStudy (campbell_st):

let \[y = (\sin(x/2)^2\] let u = sin(x/2) du/dx = 1/2cos(x/2) dy/du = 2u dy/dx = dy/du x dx/du dy/dx = 2sin(x/2) x 1/2cos(x/2) dy/dx = sin(x/2)cos(x/2) sub x = pi to get the gradient...

OpenStudy (mertsj):

When x = pi\[y'=slope=\frac{3\sin \pi}{2}=0\]

OpenStudy (campbell_st):

oops forgot the 3 at the front

OpenStudy (mertsj):

So we see that the tangent is horizontal at the given point and its equation is y =3